OpenStudy (anonymous):

will fan and give medal which pair of points has a negative slope? a.(-3,17), (2-8) b.(-3,13), (3,17) c.(4,26), (-2,-10) d.(4,14), (-2,-4)

OpenStudy (hoblos):

the slope between two points (x1,y1) and (x2,y2) is \[slope = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] you have to calculate the slope of each pair and see if it is negative

OpenStudy (anonymous):

is it b

OpenStudy (hoblos):

what is the slope for b ?

OpenStudy (hoblos):

\[\frac{ 17 - 13 }{ 3-(-3) } = \frac{ 17-13 }{ 3+3 } = \frac{ 4 }{ 6 }\] so it's positive

OpenStudy (anonymous):

what about c

OpenStudy (hoblos):

what do you get for c ?

OpenStudy (anonymous):

I don't know I just need the answer

OpenStudy (hoblos):

apply the formula i gave you above and follow the steps that i did for b

OpenStudy (campbell_st):

the easiest thing to do is plot the pairs of points... and you'll see the pair with the negative slope

OpenStudy (mrnood):

seems like the OP doesn't actually want help - just the answer.

OpenStudy (anonymous):

i do just want the answer

OpenStudy (mrnood):

@Sarahbm IF you can be bothered to accept the help given above then you will be able to do OTHER similar questions now and in the future. If you just get the answer you will know NOTHING more than you did before you started. I sincerely hope no-one posts you the straight answer

OpenStudy (campbell_st):

wow... thats harsh...

OpenStudy (anonymous):

you are right but in what life time would i you this whatto help my kids when they are older i don't care i just need the answer so i can get this work done but i do understand what you are saying.

OpenStudy (mrnood):

In the time oyu have spent waiting for an answer you could/ve used hoblos formula for the 4 pairs and done it yourself - he already did one - so it's only 3 to do

OpenStudy (hoblos):

what @MrNood said is 100% true @Sarahbm you might want to take some time and read this site's code of conduct http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. and this is not just for the person asking, it's for the person answering too, if anybody gives a direct answer he/she will be warned and would be banned if he/she insists on that. We are glad to do as much as we can to help you here, but we can't give you direct answers.

OpenStudy (anonymous):

okay im sorry i will figure it out so i am sure that it is d or a

OpenStudy (campbell_st):

here is an easy solution plot each pair or points in this online graphing calculator https://www.desmos.com/calculator you'll probably have to pull the graph down a bit then you need to know  hope it helps

OpenStudy (anonymous):

i understand it know thank you @cambell_st

OpenStudy (anonymous):

@Campbell_st

OpenStudy (campbell_st):

thats ok... I'm just saying you don't need to calculate the slope using a formula... by plotting the pairs... you see the negative slope...

OpenStudy (mrnood):

But if you DO look at the formula you will see HOW slope is calculated and it will be more obvious when you have other points to calculate. The slope is between two points is 'change in y/change in x' This is the 'word equation' being exactly the same as Hoblos' formula above. Look at the formula: y2-y1is the difference in Y between the points x2-x1 is the difference in X between the points And the slope is as given by the formula or by the word equation.

OpenStudy (campbell_st):

lol... why do extra work when a little insight and understanding eliminates numerical errors... which may occur in calculations.... and one of the fundamentals is ... answer the question...repeating a learned process doesn't show understanding... its shows the ability to repeat a process... in mho